How far will the bike in question 1 tCalculate the acceleration in each of the following cases:

(a) v = 22 ms; u = 8 ms'; t = 2.65
(b) a ball starts at rest and after 30 m its velocity has reached
4.8 ms
(c) in 15 s, a train moves 100 m from restiravel in the 10-second time period?
1
Expert's answer
2020-12-31T08:08:03-0500

(a) The acceleration is


"a=\frac{v_f-v_i}{t}=\frac{22-8}{2.65}=5.28\text{ m/s}^2."

(b) In this case, it is


"a=\frac{v^2}{2d}=\frac{4.8^2}{2\cdot30}=0.384\text{ m/s}^2."

(c) In this case, it is


"a=\frac{2d}{t^2}=\frac{2\cdot100}{15^2}=0.889\text{ m/s}^2."